# Seesaw Limit of the Nelson-Barr Mechanism

### v1

@@ I Introduction
Yet another appealing alternative is the implementation of seesaw scenarios, originally proposed to explain the smallness of neutrino masses Minkowski (1977); Yanagida (1979); Gell-Mann et al. (1979); Mohapatra and Senjanovic (1980); Schechter and Valle (1980), and later extended to charged lepton and quark masses Berezhiani (1985); Chang and Mohapatra (1987); Rajpoot (1987); Davidson and Wali (1987, 1988). Seesaw mechanisms for quarks generally occur in models where (i) new heavy quarks that mix with the SM quarks are introduced, and (ii) the conventional Yukawa interactions between the SM quarks and Higgs are forbidden by symmetries, e.g., in left-right Chang and Mohapatra (1987); Rajpoot (1987); Davidson and Wali (1987, 1988), 3-3-1 Dias et al. (2020, 2022) and $U(1)$ Jana et al. (2022); Chen and Gu (2023) extensions of the SM. The large masses of the extra quarks are responsible for suppressing the SM quark masses with respect to the electroweak scale as well as providing small mixing between the SM and the heavy quarks. The presence of extra quarks that mix with the SM fields, such as the seesaw-mediating ones, leads to rich phenomenology. If sufficiently light, they can be abundantly produced and subsequently decay at colliders due to their interactions with SM gauge bosons. Additionally, their mixing with the SM quarks induces flavour-violating interactions as well as departures from unitarity in the quark mixing, Cabibbo-Kobayashi-Maskawa (CKM), matrix. These phenomena offer multiple experimental opportunities to probe the masses and mixing of the extra quarks; for a summary of relevant constraints see Alves et al. (2024).
@@
Several connections between the strong CP and the fermion mass hierarchy problems have been explored in the literature. There are proposals with extra quarks getting massive at the PQ scale and leading to seesaw-suppressed masses for all the standard quarks Kang and Shin (1986), or tree-level (seesaw) and loop-suppressed masses for the different generations Babu and He (1989). Interrelated solutions to these problems have also been investigated in SM gauge extensions featuring the PQ mechanism, such as a 3-3-1 model Dias et al. (2018) and a non-universal $U(1)_{X}$ model Garnica et al. (2021). Furthermore, the $\bar{\theta}$ phase can be generated at the two-loop level in extended left-right models incorporating seesaw mechanisms for the SM fermions  Babu and Mohapatra (1989, 1990). On the other hand, the Froggat-Nielsen and the Peccei-Quinn mechanisms can be realised synergistically by identifying the Froggat-Nielsen flavon with the PQ axion Ema et al. (2017).
@@ II The Nelson-Barr mechanism
The contributions at tree level to $\bar{\theta}$ coming from the quark sector ($\theta_{q}$) vanish because $\arg(\det\mathcal{M}^{d+B})=\arg(\det m^{u})=0$. Moreover, since the $CP$ symmetry is only broken (spontaneously) through $\mathcal{M}^{Bd}$, the bare $\theta_{QCD}$ term also vanishes. Consequently, the effective parameter $\bar{\theta}=\theta_{QCD}+\theta_{q}$ vanishes at tree level.
@@ III The seesaw limit of the Nelson-Barr mechanism
where $\hat{Y}^{d}=\mathrm{diag}(y_{d},y_{s},y_{b})$ is the diagonal Yukawa coupling matrix for the down-type quarks in the SM, and $U$ is a generic $3\times 3$ unitary matrix. This inversion is particularly simple in the seesaw limit; away from this limit a more complicated parametrization similar to the ones in Refs. Cherchiglia and Nishi (2020); Alves et al. (2023) will be needed. Without loss of generality, we can choose a basis in which $\mathcal{M}^{B}$ is diagonal and $\mathcal{Y}^{d}=O_{d_{L}}\hat{\mathcal{Y}}^{d}$, with orthogonal $O_{d_{L}}$ and diagonal $\hat{\mathcal{Y}}^{d}$. Finally, we can see that if $\mathcal{Y}^{d}$ has order-one entries, $\mathcal{M}^{Bd}$ is larger than $\mathcal{M}^{B}$ by, at least, $1/y_{b}\sim 10^{2}$. This is clear if we substitute Eq. (18) into Eq. (17), obtaining
@@
where $V_{0}$ is, e.g., the standard parametrization. These additional phases arise because the Yukawa $Y^{d}$ in Eq. (6) loses rephasing freedom from the left to keep $\mathcal{Y}^{d}$ real.
@@
In such cases, instead of Eq. (18), we can use the formulae derived in Refs. Cherchiglia and Nishi (2020); Alves et al. (2023) to find benchmarks satisfying the current quark mixing picture. The exact parametrization involves solving the first equation of (11) for $\mathcal{Y}^{d}$. To that end, we first rewrite it as
@@ IV Nelson-Barr seesaw for the down-type quarks
In the second case, the singular values $\hat{\mathcal{Y}}^{d}_{ii}$ follow the hierarchy of the SM Yukawas $y_{i}=(y_{d},y_{s},y_{b})$ allowing for a variation of 25%. We can see in the figure in the generic $\mathcal{Y}^{d}$ set that $\mathcal{Y}^{d}$ of the same order requires hierarchical $\epsilon$ to account for the hierarchy of the SM Yukawa $Y^{d}$. On the other hand, as $\mathcal{Y}^{d}$ becomes hierarchical, a less hierarchical $\epsilon$ is necessary. In the extreme case of $\mathcal{Y}^{d}$ following the hierarchy of the SM Yukawa (hierarchical $\mathcal{Y}^{d}$ set), non-hierarchical values for $\epsilon$ are allowed.
Figure 1: Singular values $\lambda_{i}^{X}$ of $X=\mathcal{Y}^{d}$ as a function of singular values for $X=\epsilon$. The filled (hollow) points denote the generic (hierarchical) $\mathcal{Y}^{d}$ case described in Eq. (31). Diagonal lines represent $\lambda_{i}^{\mathcal{Y}^{d}}\lambda_{i}^{\epsilon}=y_{i}$ (no sum), where $y_{i}$ are the SM Yukawa couplings.
@@
Finally, we show in Fig. 3 the ratio of the heavy masses in Eq. (21) compared to their lightest mass as a function of $|\epsilon^{-1}|$. In accordance with Eq. (22), we see that larger $|\epsilon^{-1}|$ tends to lead to a larger hierarchy among the heavy quarks. For the hierarchical $\mathcal{Y}^{d}$ set where the entries of $\epsilon$ are all of the same order, the heavy masses are also non-hierarchical.
@@ V Nelson-Barr seesaw for the up-type quarks
Figure 4: Singular values $\lambda_{i}^{\mathcal{Y}^{u}}$ as a function of singular values $\lambda_{i}^{\epsilon}$. The hollow points comply with $\lambda_{1}^{\epsilon}/\lambda_{2}^{\epsilon}>0.5$. The orange (blue) line corresponds to the ratio between the up (charm) Yukawa to the top Yukawa, in the SM.
In figure 4, we show the ratio among the two smallest singular values of $\mathcal{Y}^{u}$ to the third, against $\lambda^{\epsilon}_{i}$. As the ratio $\lambda_{1}^{y^{u}}/\lambda_{3}^{y^{u}}$ approaches the SM value (orange line), the singular values of $\epsilon$ are at most one order of magnitude apart. In this regime, the observed hierarchy among the first and second family masses comes from the hierarchical structure of $\mathcal{Y}^{u}$. On the other hand, as the singular values of $\epsilon$ differ by roughly 3 orders of magnitude (which is the observed hierarchy among the first and second up-quark family in the SM), the ratios $\lambda_{1}^{y^{u}}/\lambda_{3}^{y^{u}}$, $\lambda_{2}^{y^{u}}/\lambda_{3}^{y^{u}}$ approach similar values. More precisely, they are close to the blue line, which represents the ratio $m_{c}/m_{t}$. In this case, the observed hierarchy among up and charm quark masses resides in $\epsilon$. In the figure, the hollow points comply with $\lambda_{1}^{\epsilon}/\lambda_{2}^{\epsilon}>0.5$.
@@
Figure 6: The ratio $M^{T}_{2}/M^{T}_{1}$ of the heavy quarks as a function of the norm of $\epsilon^{-1}$ in Eq. (33).
@@ VI Conclusions
Although the SM is very successful in explaining a plethora of phenomena, there are still some puzzling issues, such as the strong CP problem and the observed fermion mass hierarchies, to cite a few. In this work, seeking a common solution to these problems, we have identified and investigated the “seesaw limit” of the Nelson-Barr mechanism. We have studied in detail how the introduction of three heavy VLQs, that mix with the Standard Model fields, can provide a natural mechanism to suppress the SM down quark masses in relation to the EW scale. In particular, we have seen that the source of CP violation in the quark sector, i.e., the SM-VLQ-mixing term $\mathcal{M}^{Bd}$ in Eqs. (2) and (3), is also behind the seesaw suppression of light quark masses. Since these VLQs are connected to the Nelson-Barr mechanism, not only the strong CP problem can be addressed, but also the couplings between them and the standard quarks are typically hierarchical. This in turn allows our model to naturally evade strong bounds from flavour observables coming from first and second quark families. We have also shown in a similar model with two up-type VLQs that the hierarchy among the SM up quarks can be partly explained by the hierarchy of the small seesaw parameters.

### v2

@@ I Introduction
Yet another appealing alternative is the implementation of seesaw scenarios, originally proposed to explain the smallness of neutrino masses Minkowski (1977); Yanagida (1979); Gell-Mann et al. (1979); Mohapatra and Senjanovic (1980); Schechter and Valle (1980), and later extended to charged lepton and quark masses Berezhiani (1985); Chang and Mohapatra (1987); Rajpoot (1987); Davidson and Wali (1987, 1988). Seesaw mechanisms for quarks generally occur in models where (i) new heavy quarks that mix with the SM quarks are introduced, and (ii) the conventional Yukawa interactions between the SM quarks and Higgs are forbidden by symmetries, e.g., in left-right Chang and Mohapatra (1987); Rajpoot (1987); Davidson and Wali (1987, 1988), 3-3-1 Dias et al. (2020, 2022) and $U(1)$ Jana et al. (2022); Mohanta and Patel (2022); Chen and Gu (2023) extensions of the SM. The large masses of the extra quarks are responsible for suppressing the SM quark masses with respect to the electroweak scale as well as providing small mixing between the SM and the heavy quarks. The presence of extra quarks that mix with the SM fields, such as the seesaw-mediating ones, leads to rich phenomenology. If sufficiently light, they can be abundantly produced and subsequently decay at colliders due to their interactions with SM gauge bosons. Additionally, their mixing with the SM quarks induces flavour-violating interactions as well as departures from unitarity in the quark mixing, Cabibbo-Kobayashi-Maskawa (CKM), matrix. These phenomena offer multiple experimental opportunities to probe the masses and mixing of the extra quarks; for a summary of relevant constraints see Alves et al. (2024).
@@
Several connections between the strong CP and the fermion mass hierarchy problems have been explored in the literature. There are proposals with extra quarks getting massive at the PQ scale and leading to seesaw-suppressed masses for all the standard quarks Kang and Shin (1986), or tree-level (seesaw) and loop-suppressed masses for the different generations Babu and He (1989). Interrelated solutions to these problems have also been investigated in SM gauge extensions featuring the PQ mechanism, such as a $SU(3)_{H}$ extension Berezhiani and Khlopov (1990), a 3-3-1 model Dias et al. (2018) and a non-universal $U(1)_{X}$ model Garnica et al. (2021). Furthermore, the $\bar{\theta}$ phase can be generated at the two-loop level in extended left-right models incorporating seesaw mechanisms for the SM fermions  Babu and Mohapatra (1989, 1990). On the other hand, the Froggat-Nielsen and the Peccei-Quinn mechanisms can be realised synergistically by identifying the Froggat-Nielsen flavon with the PQ axion Ema et al. (2017); Calibbi et al. (2017).
@@ II The Nelson-Barr mechanism
The contributions at tree level to $\bar{\theta}$ coming from the quark sector ($\theta_{q}$) vanish because $\arg(\det\mathcal{M}^{d+B})=\arg(\det m^{u})=0$. Moreover, since the $CP$ symmetry is only broken (spontaneously) through $\mathcal{M}^{Bd}$, the bare $\theta_{QCD}$ term also vanishes. Consequently, the effective parameter $\bar{\theta}=\theta_{QCD}+\theta_{q}$ vanishes at tree level. The leading one-loop corrections can also be made small by, e.g., suppressing the Yukawa couplings of the VLQs with the CP-breaking scalars. There are, however, irreducible contributions at 3-loops Valenti and Vecchi (2021b) which might be more important. For generic points, these contributions are relevant for two or more up-type VLQs but they are still compatible with current bounds Alves et al. (2023). Their importance in the seesaw limit of Sec. III is discussed in Secs. IV and V.
@@ III The seesaw limit of the Nelson-Barr mechanism
where $\hat{Y}^{d}=\mathrm{diag}(y_{d},y_{s},y_{b})$ is the diagonal Yukawa coupling matrix for the down-type quarks in the SM, and $U$ is a generic $3\times 3$ unitary matrix. This relation has been obtained by assuming that $\mathcal{H}^{d}$ is diagonalised by the CKM matrix, i.e. $V_{CKM}^{\dagger}\mathcal{H}_{d}V_{CKM}\simeq\mathrm{diag}(m_{d}^{2},m_{s}^{2},m_{b}^{2})$; therefore, all points generated from Eq. (18) are automatically in agreement with the observed values of the CKM matrix parameters. This inversion is particularly simple in the seesaw limit; away from this limit a more complicated parametrization similar to the ones in Refs. Cherchiglia and Nishi (2020); Alves et al. (2023) will be needed. Without loss of generality, we can choose a basis in which $\mathcal{M}^{B}$ is diagonal and $\mathcal{Y}^{d}=O_{d_{L}}\hat{\mathcal{Y}}^{d}$, with orthogonal $O_{d_{L}}$ and diagonal $\hat{\mathcal{Y}}^{d}$. Finally, we can see that if $\mathcal{Y}^{d}$ has order-one entries, $\mathcal{M}^{Bd}$ is larger than $\mathcal{M}^{B}$ by, at least, $1/y_{b}\sim 10^{2}$. This is clear if we substitute Eq. (18) into Eq. (17), obtaining
@@
where $V_{0}$ is, e.g., the standard parametrization. These additional phases arise because the Yukawa $Y^{d}$ in Eq. (6) loses rephasing freedom from the left to keep $\mathcal{Y}^{d}$ real. Notice that, in the particular case that the generic matrices $U$ and $O_{d_{L}}$ are diagonal, only $\mathcal{M}^{Bd}$ and $\epsilon$ are nondiagonal owing to the CKM matrix. As in the SM, we do not aim to explain the origin of the hierarchy exhibited by the CKM but rather provide a solution for the hierarchy among the quark masses. In other words, the entries of $\hat{\mathcal{Y}}^{d}$ can be much larger than the down Yukawa matrix in the SM, while their mixing comes from phenomenological observation.
@@
In such cases, instead of Eq. (18), we can use the formulae derived in Refs. Cherchiglia and Nishi (2020); Alves et al. (2023) to find benchmarks satisfying the current quark mixing picture. In particular, the CKM matrix is automatically taken into account. The exact parametrization involves solving the first equation of (11) for $\mathcal{Y}^{d}$. To that end, we first rewrite it as
@@ IV Nelson-Barr seesaw for the down-type quarks
In the second case, the singular values $\hat{\mathcal{Y}}^{d}_{ii}$ follow the hierarchy of the SM Yukawas $y_{i}=(y_{d},y_{s},y_{b})$ allowing for a variation of 25%. We can see in the figure in the generic $\mathcal{Y}^{d}$ set that $\mathcal{Y}^{d}$ of the same order requires hierarchical $\epsilon$ to account for the hierarchy of the SM Yukawa $Y^{d}$. On the other hand, as $\mathcal{Y}^{d}$ becomes hierarchical, a less hierarchical $\epsilon$ is necessary. In the extreme case of $\mathcal{Y}^{d}$ following the hierarchy of the SM Yukawa (hierarchical $\mathcal{Y}^{d}$ set), non-hierarchical values for $\epsilon$ are allowed. Finally, one could wonder if the observed features in Fig. 1 could be due to some hidden hierarchy in the matrices $U$ and $O_{d_{L}}$. We have checked that, even in the case that they are equal to the identity matrix, the overall behaviour is maintained, i.e., the hierarchies of the down Yukawas can be attributed to $\mathcal{Y}^{d}$, to $\epsilon$ or to an interplay between them.
[…]
@@
To get a sense of the spectrum for the VLQs, we show in Fig. 3 the ratio of the heavy masses in Eq. (21) compared to their lightest mass as a function of $|\epsilon^{-1}|$. In accordance with Eq. (22), we see that larger $|\epsilon^{-1}|$ tends to lead to a larger hierarchy among the heavy quarks. For the hierarchical $\mathcal{Y}^{d}$ set where the entries of $\epsilon$ are all of the same order, the heavy masses are also non-hierarchical.
@@
Finally, we show in Fig. 4 the estimate to the 3-loop irreducible contribution to $\bar{\theta}$ coming from VLQ exchange against $|\epsilon^{-1}|$. The dominant contribution is expected to be Valenti and Vecchi (2021a)
[…]
with model dependent order one coefficient, and we define $F(A)=A/\operatorname{Tr}[A]$. The distinction among hollow, filled and dark (opaque) points follows the convention of Fig. 1. We, however, have already filtered out the points that violate the perturbativity of the Yukawas. We can see that the dark points, corresponding to the case where $\mathcal{Y}^{d}\propto{\mathbbm{1}}_{3}$ carries no hierarchy, are severely excluded by the limits on $\bar{\theta}$ from the neutron electric dipole moment. The other subsets, although severely constrained, contain points that pass the limit. In particular, the case of hollow points, cf. (31), where $\mathcal{Y}^{d}$ carries the hierarchy among the SM Yukawas, contains a reasonable quantity of viable points.
Figure 4: Estimate (35) of the irreducible 3-loop contributions to $\bar{\theta}$ against $|\epsilon^{-1}|$ in Eq. (33) for three down-type VLQs in the seesaw limit. The differences among hollow, filled and dark points follow the convention of Fig. 1.
@@ V Nelson-Barr seesaw for the up-type quarks
Figure 5: Singular values $\lambda_{i}^{\mathcal{Y}^{u}}$ as a function of singular values $\lambda_{i}^{\epsilon}$. The hollow points comply with $\lambda_{1}^{\epsilon}/\lambda_{2}^{\epsilon}>0.5$, while for the darker points we require $\lambda_{1}^{\mathcal{Y}^{u}}/\lambda_{2}^{\mathcal{Y}^{u}}>0.6$. The orange (blue) line corresponds to the ratio between the up (charm) Yukawa to the top Yukawa, in the SM.
In figure 5, we show the ratio among the two smallest singular values of $\mathcal{Y}^{u}$ to the third, against $\lambda^{\epsilon}_{i}$. As the ratio $\lambda_{1}^{y^{u}}/\lambda_{3}^{y^{u}}$ approaches the SM value (orange line), the singular values of $\epsilon$ are at most one order of magnitude apart. In this regime, the observed hierarchy among the first and second family masses comes from the hierarchical structure of $\mathcal{Y}^{u}$. This feature can be more clearly seen when considering the hollow points, which comply with $\lambda_{1}^{\epsilon}/\lambda_{2}^{\epsilon}>0.5$. On the other hand, as the singular values of $\epsilon$ differ by roughly 3 orders of magnitude (which is the observed hierarchy among the first and second up-quark family in the SM), the ratios $\lambda_{1}^{y^{u}}/\lambda_{3}^{y^{u}}$, $\lambda_{2}^{y^{u}}/\lambda_{3}^{y^{u}}$ approach similar values. More precisely, they are close to the blue line, which represents the ratio $m_{c}/m_{t}$. In this case, the observed hierarchy among up and charm quark masses resides in $\epsilon$. This feature can be seen for the darker points, selected with $\lambda_{1}^{\mathcal{Y}^{u}}/\lambda_{2}^{\mathcal{Y}^{u}}>0.6$, where $\lambda_{1}^{\epsilon}$ and $\lambda_{2}^{\epsilon}$ are at least two orders of magnitude apart.
@@
Figure 7: The ratio $M_{2}^{T}/M_{1}^{T}$ of the heavy quarks as a function of the norm of $\epsilon^{-1}$ in Eq. (33).
Finally, in Fig. 8 we show the estimate to the irreducible contribution to $\bar{\theta}$ at 3-loops coming from the exchange of the VLQs against $|\epsilon^{-1}|$. The following flavor invariant is expected to dominate for two or more up-type VLQs Valenti and Vecchi (2021a):
[…]
where we are adopting the basis where $Y^{T}=\tilde{Y}^{T}$ in the basis where $Y^{u}=\hat{Y}^{u}$ is diagonal, and the function $F$ is defined below (35). This estimation is expected to be corrected by order one prefactors depending on the model. We compare generic points in blue, already quantified in Ref. Alves et al. (2023), with the points in the seesaw limit where the singular values of $\mathcal{M}^{Tu}$ is larger than the singular values of $\mathcal{M}^{T}$ by at least one order of magnitude. We can see that plenty of points still obey $\bar{\theta}\lesssim 10^{-10}$ even if with a possible enhancement of an order of magnitude. As for the 3 VLQs case, we have filtered out the points that violate perturbativity.
Figure 8: Estimate (36) of the irreducible 3-loop contributions to $\bar{\theta}$ against $|\epsilon^{-1}|$ in Eq. (33) for two up-type VLQs. Points in black satisfy the seesaw limit condition while those in blue represent generic parameter choices.
@@ VI Conclusions
Although the SM is very successful in explaining a plethora of phenomena, there are still some puzzling issues, such as the strong CP problem and the observed fermion mass hierarchies, to cite a few. In this work, seeking a common solution to these problems, we have identified and investigated the “seesaw limit” of the Nelson-Barr mechanism. We have studied in detail how the introduction of three heavy VLQs, that mix with the Standard Model fields, can provide a natural mechanism to suppress the SM down quark masses in relation to the EW scale. In particular, we have seen that the source of CP violation in the quark sector, i.e., the SM-VLQ-mixing term $\mathcal{M}^{Bd}$ in Eqs. (2) and (3), is also behind the seesaw suppression of light quark masses. Since these VLQs are connected to the Nelson-Barr mechanism, not only the strong CP problem can be addressed, but also the couplings between them and the standard quarks are typically hierarchical. This in turn allows our model to naturally evade strong bounds from flavour observables coming from first and second quark families. We have also shown in a similar model with two up-type VLQs that the hierarchy among the SM up quarks can be partly explained by the hierarchy of the small seesaw parameters. For both cases, the irreducible 3-loop contributions to $\bar{\theta}$ severely restrict the parameter space of the model, especially for the case of three down-type VLQs. For the latter, such a restriction is weaker for the cases where the CP conserving Yukawa couplings $\mathcal{Y}^{d}$ of the down quarks in the CP basis are mostly responsible for the mass hierarchies.